In how many years will Mr. Sison be 7 times that of his son if he is now 69 years old and his son is 3?
Please give me he representation of your answer thank you!!!!

Answer by checkley77(12844)

69+x=7(3+x)
69+x=21+7x
x-7x=21-69
-6x=-48
x=-48/-6
x=8 years from now dad will be 7 times as old as his son.
Proof:
69+8=8(3+8)
77=7*11
77=77